### v3

# Classification of high-dimensional data with spiked covariance matrix structure

Abstract
We study the classification problem for high-dimensional data with $n$ observations on $p$ features where the $p\times p$ covariance matrix $\Sigma$ exhibits a spiked eigenvalue structure and the vector $\zeta$, given by the difference between the whitened mean vectors, is sparse. We analyze an adaptive classifier (adaptive with respect to the sparsity $s$) that first performs dimension reduction on the feature vectors prior to classification in the dimensionally reduced space, i.e., the classifier whitens the data, then screens the features by keeping only those corresponding to the $s$ largest coordinates of $\zeta$ and finally applies Fisher linear discriminant on the selected features. Leveraging recent results on entrywise matrix perturbation bounds for covariance matrices, we show that the resulting classifier is Bayes optimal whenever $n\rightarrow\infty$ and $s\sqrt{n^{-1}\ln p}\rightarrow 0$. Notably, our theory also guarantees Bayes optimality for the corresponding quadratic discriminant analysis (QDA). Experimental results on real and synthetic data further indicate that the proposed approach is competitive with state-of-the-art methods while operating on a substantially lower-dimensional representation.
1 Introduction
Classification is one of the most important and widely studied inference tasks in statistics and machine learning. Among standard classifiers, the Fisher linear discriminant analysis (LDA) rule is especially popular for its ease of implementation and interpretation. More specifically, suppose that we are given a $p$-variate random vector $\mathbf{Z}$ drawn from a mixture of two multivariate normal distributions $\pi_{1}\mathcal{N}_{p}(\mu_{1},\Sigma)+(1-\pi_{1})\mathcal{N}_{p}(\mu_{2},\Sigma)$ and our goal is to classify $\mathbf{Z}$ into one of the two classes. The Fisher LDA rule is then given by
$\Upsilon_{F}(\mathbf{Z})=\begin{cases}1&\text{if $\bigl(\Sigma^{-1}(\mu_{2}-\mu_{1})\bigr)^{\top}\bigl(\mathbf{Z}-\tfrac{\mu_{1}+\mu_{2}}{2}\bigr)\leq\ln\tfrac{\pi_{1}}{1-\pi_{1}}$,}\\ 2&\text{if $\bigl(\Sigma^{-1}(\mu_{2}-\mu_{1})\bigr)^{\top}\bigl(\mathbf{Z}-\tfrac{\mu_{1}+\mu_{2}}{2}\bigr)>\ln\tfrac{\pi_{1}}{1-\pi_{1}}$,}\end{cases}$ (1.1)
provided that $\Sigma$ is positive definite. The Fisher rule $\Upsilon_{F}$ is the Bayes decision rule, i.e., it achieves the smallest mis-classification error with respect to 0-11 loss for classifying $\mathbf{Z}$. $\Upsilon_{F}$ is, however, not directly applicable in practice as it involves the unknown parameters $\Sigma^{-1}$, $\mu_{1}$ and $\mu_{2}$; we thus usually compute the sample covariance matrix $\hat{\Sigma}$ and the sample means $\bar{X}_{1}$ and $\bar{X}_{2}$ from a given training data set with $n$ observations and then plugged these quantities into Eq.˜1.1; the resulting classifier is termed as the plug-in LDA rule $\hat{\Upsilon}_{F}$.
The classification accuracy of $\hat{\Upsilon}_{F}$ is well-understood in the low-dimensional regime where $p\ll n$. In particular $\hat{\Upsilon}_{F}$ is asymptotically optimal, i.e., it achieves the Bayes error rate as $n\rightarrow\infty$ for fixed $p$. This behavior, however, might no longer holds in the high-dimensional settings $n\asymp p$ or even $n\ll p$ where the sample covariance matrix $\hat{\Sigma}$ is singular. Indeed, Bickel & Levina (2004) proved that the error rate for $\hat{\Upsilon}_{F}$ where we replace $\Sigma^{-1}$ with the Moore–Penrose pseudo-inverse of $\hat{\Sigma}$ could be as bad as random guessing and that, furthermore, the naive Bayes (NB) rule which ignores the correlation structure in $\Sigma$ typically outperforms $\hat{\Upsilon}_{F}$.
Continuing this line of inquiry, other independence rules (IR) have been proposed in Tibshirani et al. (2002) and Fan & Fan (2008) but these classifiers have two potentially major drawbacks, namely (1) the accumulation of error in estimating $\mu_{1}$ and $\mu_{2}$ and (2) mis-specification of the covariance matrix $\Sigma$. In particular Fan & Fan (2008) demonstrated that the noise accumulation in estimating $\mu_{1}$ and $\mu_{2}$ alone is sufficient to degrade the performance of IR classifiers, Shao et al. (2011) extended this result to show that the noise accumulation can also lead to the plug-in LDA classifier being, asymptotically, no better than random guessing even when $\Sigma$ is known, and finally Fan et al. (2012) showed that ignoring the correlations structure in $\Sigma$ prevents IR classifiers from achieving Bayes optimality. These results indicate the need for imposing sparsity conditions on the features and the important role of feature selection for mitigating the estimation errors associated with growing dimensionality.
To address the above limitations of IR classifiers, Shao et al. (2011) proposed thresholding of both $\Sigma$ and $\mu_{2}-\mu_{1}$; this is similar to the motivation for regularized covariance estimators in Bickel & Levina (2009). In contrast, Cai & Liu (2011); Fan et al. (2012); Mai et al. (2012); Witten & Tibshirani (2011) and Cai & Zhang (2019) imposed sparsity conditions on the discriminant direction $\beta=\Sigma^{-1}(\mu_{2}-\mu_{1})$ and use penalized estimation approaches to recover $\beta$. Cai & Liu (2011) also noted two potential advantages to this approach, namely that the assumption of sparsity on $\beta$ is less restrictive than assuming sparsity for both $\Sigma$ (or $\Sigma^{-1}$) and $\mu_{2}-\mu_{1}$, and secondly $\Upsilon_{F}$ only depends on $\Sigma^{-1}$ and $\mu_{2}-\mu_{1}$ through their product $\Sigma^{-1}(\mu_{2}-\mu_{1})$ and thus consistent estimation of $\beta$ is sufficient. The sparsity assumption on $\beta$ leads to procedures and results that resemble those for high-dimensional linear regression even though the classification problem is generally not formulated in terms of a linear model. For example, Cai & Liu (2011) and Cai & Zhang (2019) considered a linear programming approach similar to the Dantzig selector (Candes & Tao, 2005) while Mai et al. (2012) studied a sparse discriminant analysis rule that used the Lasso (Tibshirani, 1996). Nevertheless it had been observed that, empirically, these approaches can lead to classification rules which select a larger number of features than necessary, and one possible explanation is that the correlation structure in $\Sigma$ also induced correlations among the entries for any estimate $\hat{\beta}$ of $\beta$.
In this paper we consider a different approach where we first perform dimension reduction on the feature vectors (using PCA) prior to classification (using LDA) in the dimensionally reduced space. As PCA is an important and ubiquitous pre-processing step in high-dimensional data analysis, there is a sizable number of work devoted to this approach. We refer to a generic classifier from this combination as $\mathrm{lda}\circ\mathrm{pca}$. For example Section 9.1 of Jolliffe (2002) provides a detailed review of combining LDA with different variants of PCA in the low-dimensional setting while Niu et al. (2015) proposed the use of the reduced rank LDA together with class-conditional PCA in the high-dimensional settings. In terms of applications, $\mathrm{lda}\circ\mathrm{pca}$ is also used for faces and images recognition (Zhao et al., 1998; Belhumeur et al., 1997; Prasad et al., 2010), and recovering genetic patterns (Jombart et al., 2010)
While $\mathrm{lda}\circ\mathrm{pca}$ classifiers arise quite naturally, their theoretical properties in the high-dimensional setting remains an open problem. In particular their analysis requires possibly different techniques and assumptions compared to those based on direct estimation of the discriminant direction $\beta$. More specifically consistency results for $\beta$ are usually based on ideas from high-dimensional regression including assumptions on bounded and concentrated eigenvalues of $\Sigma$, see e.g., Bickel & Levina (2004); Fan & Fan (2008); Shao et al. (2011); Cai & Zhang (2019). However Wang & Fan (2017) noted that bounded eigenvalues are incompatible with the presence of strong signals (eigenvalues) in the data and might be problematic in fields such as genomics, economics and finance. These assumptions are nevertheless imposed mainly due to the limitation of quantifying the estimation error for $\Sigma$ in terms of the spectral norm difference for $\hat{\Sigma}-\Sigma$. In contrast the idealized setting for dimension reduction via PCA is when $\Sigma$ contains a small subset of signal eigenvalues that accounts for most of the variability in $\Sigma$. This idealized setting for PCA is also distinct from the idealized setting for graphical models wherein $\Sigma^{-1}$ is typically assumed to be sparse.
In summary, our contributions in this paper are as follows. We analyze the theoretical properties of a prototypical $\mathrm{lda}\circ\mathrm{pca}$ classifier under a spiked-covariance structure assumption – a widely-adopted covariance model for high-dimensional data – where $\Sigma$ contains a few large eigenvalues that are well-separated from the remaining (small) eigenvalues. In particular we show in Section 3 that $\mathrm{lda}\circ\mathrm{pca}$ is asymptotically Bayes-optimal as $n\rightarrow\infty$ and $n^{-1}\ln p\rightarrow 0$. This is, to the best of our knowledge, the first Bayes optimal consistency result for classification after performing dimension reduction via PCA. In Section 4 we demonstrate empirically, for both simulated and real data, that $\mathrm{lda}\circ\mathrm{pca}$ selects fewer features while also having error rates that are competitive with existing classifiers based on estimating the discriminant direction $\beta$. The theoretical and numerical results provide a clear example of the synergy linking dimension reduction with classification. Finally, in Section 5 we extend $\mathrm{lda}\circ\mathrm{pca}$ to classify data with (1) $K\geq 3$ classes or (2) two classes but with unequal covariance matrices or (3) feature vectors that are elliptical distributed but not necessarily multivariate normal.
2 Methodology
2.1 Notation and settings
For a vector $x\in\mathbb{R}^{p}$, the conventional $\ell_{0}$ quasi-norm and the $\ell_{1}$, $\ell_{2}$ and $\ell_{\infty}$ norms are denoted by $\left\lVert x\right\rVert_{0}$, $\left\lVert x\right\rVert_{1}$, $\left\lVert x\right\rVert_{2}$ and $\left\lVert x\right\rVert_{\infty}$, respectively. For $p\in\mathbb{N}$, we denote the set $\{1,...,p\}$ by $[p]$. Given $x\in\mathbb{R}^{p}$ and a non-empty set $\mathcal{A}\subset[p]$, we write $x_{\mathcal{A}}=(x_{j},j\in\mathcal{A})$ to denote the column vector obtained by keeping only the elements in $x$ whose indices belong to $\mathcal{A}$. The operation $\circ$’, when applied to matrices, represents the Hadamard (entrywise) product. For $i\in[p]$, $\mathbf{e}_{i}^{(p)}$ is the $i$th standard basis vector of $\mathbb{R}^{p}$ and $\bm{1}_{p}$ is a vector in $\mathbb{R}^{p}$ whose elements are all $1$; we also write $e_{i}$ and $\bm{1}$ when the choice of $p$ is clear from context and $\mathcal{I}_{p}$ stands for the identity matrix in $\mathbb{R}^{p\times p}$. For $x,y\in\mathbb{R}^{p}$, the standard Euclidean inner product between $x$ and $y$ is denoted as $\langle x,y\rangle\vcentcolon=y^{\top}x$. For a matrix $M\in\mathbb{R}^{p\times q}$, the Frobenius norm and spectral norm of $M$ are written as $\left\lVert M\right\rVert_{F}$ and $\left\lVert M\right\rVert_{2}$, respectively. We will omit the subscript in $\left\lVert\cdot\right\rVert_{2}$ when it is clear from context that the spectral norm or $\ell_{2}$ norm is intended.
The two-to-infinity norm of $M$ is defined as
$\left\lVert M\right\rVert_{2\rightarrow\infty}\vcentcolon=\sup_{\left\lVert x\right\rVert_{2}=1}\left\lVert Mx\right\rVert_{\infty}\equiv\max_{i\in[p]}\|M_{i}\|_{2}$ (2.1)
[…]
In the subsequent discussion we shall generally assume, unless specified otherwise, that we have access to training set $\{X_{11},\dots,X_{1n_{1}}\}$ and $\{X_{21},\dots,X_{2n_{2}}\}$ whose elements are independently and identically distributed random vector from the $p$-variate distributions $\mathcal{N}_{p}(\mu_{1},\Sigma)$ (class 1) and $\mathcal{N}_{p}(\mu_{2},\Sigma)$ (class 2). The sample means for each class and the pooled sample covariance matrix are denoted as
$\bar{X}_{1}=\frac{1}{n_{1}}\sum_{j=1}^{n_{1}}X_{1j},\qquad\bar{X}_{2}=\frac{1}{n_{2}}\sum_{j=1}^{n_{2}}X_{2j},$ (2.2)
$\hat{\Sigma}=\frac{1}{n}\Bigl(\sum_{j=1}^{n_{1}}(X_{1j}-\bar{X}_{1})(X_{1j}-\bar{X}_{1})^{\top}+\sum_{j=1}^{n_{2}}(X_{2j}-\bar{X}_{2})(X_{2j}-\bar{X}_{2})^{\top}\Bigr)$ (2.3)
On numerous occasions we also need a variant of the sample covariance matrix where we replace the sample means with the true means. We denote this matrix as
$\hat{\Sigma}_{0}=\frac{1}{n}\Bigl(\sum_{j=1}^{n_{1}}(X_{1j}-\mu_{1})(X_{1j}-\mu_{1})^{\top}+\sum_{j=1}^{n_{2}}(X_{2j}-\mu_{2})(X_{2j}-\mu_{2})^{\top}\Bigr).$ (2.4)
2.2 Linear discriminant analysis and whitening matrix
Let $\Sigma$ be a $p\times p$ positive definite matrix. The whitening matrix $\mathcal{W}$ is a linear transformation satisfying $\mathcal{W}\mathcal{W}^{\top}=\Sigma^{-1}$, and is generally used to decorrelate random variables and scale their variances to $1$. The whitening transformation is unique only up to orthogonal transformation as $\mathcal{W}TT^{\top}\mathcal{W}^{\top}=\mathcal{W}\mathcal{W}$ for any $p\times p$ orthogonal matrix $T$; see Kessy et al. (2015) for a comparison between several common choices of whitening transformation. Hence, for this paper, we take $\mathcal{W}=\Sigma^{-1/2}$ as the unique positive semidefinite square root of $\Sigma^{-1}$. Note that $\Sigma^{-1/2}$ is the whitening transformation which minimizes the expected mean square error between the original data and the whitened data (Eldar & Oppenheim, 2003). Given $\mathcal{W}$, we define the whitened direction as $\zeta=\mathcal{W}(\mu_{2}-\mu_{1})$ and let $\mathcal{S}_{\zeta}=\{j:\zeta_{j}\neq 0\}$. We referred to the elements of $\mathcal{S}_{\zeta}$ as the whitened coordinates or variables. The Fisher linear discriminant rule is then equivalent to
$\Upsilon_{F}(\mathbf{Z})=\begin{cases}1&\text{if $\zeta^{\top}\bigl(\mathcal{W}\bigl(\mathbf{Z}-\tfrac{\mu_{1}+\mu_{2}}{2}\bigr)\bigr)=\zeta_{\mathcal{S}_{\zeta}}^{\top}\bigl(\mathcal{W}\bigl(\mathbf{Z}-\tfrac{\mu_{1}+\mu_{2}}{2}\bigr)\bigr)_{\mathcal{S}_{\zeta}}\leq\ln\tfrac{\pi_{1}}{1-\pi_{1}}$},\\ 2&\text{otherwise},\end{cases}$ (2.5)
where $\mathbf{Z}\sim\pi_{1}\mathcal{N}_{p}(\mu_{1},\Sigma)+(1-\pi_{1})\mathcal{N}_{p}(\mu_{2},\Sigma)$. Recall that, for a vector $\xi$ and a set of indices $\mathcal{A}$, the vector $\xi_{\mathcal{A}}$ is obtained from $\xi$ by keeping only the elements indexed by $\mathcal{A}$. Since $\beta=\Sigma^{-1}(\mu_{2}-\mu_{1})$ is the Bayes direction, a significant portion of previous research is devoted to recovering $\beta$ and the discriminative set $\mathcal{S}_{\beta}=\{j:\beta_{j}\neq 0\}$. We refer to the elements of $\mathcal{S}_{\beta}$ as the discriminative coordinates or variables. We emphasize that in general the discriminative set and whitened set are not the same, except for when $\Sigma$ has certain special structures, e.g., $\Sigma$ being diagonal.
If $p\ll n$ then the empirical whitening matrix $\hat{\mathcal{W}}=\hat{\Sigma}^{-1/2}$ is well-defined and the empirical whitened variables $\hat{\zeta}=\hat{\mathcal{W}}(\bar{X}_{2}-\bar{X}_{1})$ are approximately decorrelated in the plug-in LDA rule, i.e., by the law of large numbers, $\hat{\mathcal{W}}\Sigma\hat{\mathcal{W}}\rightarrow\mathcal{I}_{p}$ and hence $\mathrm{Var}[\hat{\zeta}]\approx c\mathcal{I}_{p}$ for some constant $c>0$. In contrast, the coordinates of the estimated discriminant direction $\hat{\beta}$ are not decorrelated since $\mathrm{Var}[\hat{\beta}]\approx c\Sigma^{-1}$. It is thus easier to quantify the impact of any arbitrary whitened feature toward the classification accuracy than to quantify the impact of an arbitrary raw feature.
2.3 Combining LDA and PCA
Eq.˜2.5 can be viewed under the framework of first performing dimension reduction and then doing classification in the lower-dimensional spaces with the main focus being that thresholding of the discriminant direction corresponds to dropping a subset of the raw features while thresholding of the whitened direction corresponds to dropping a subset of the whitened features. We now discuss the estimation of the whitening matrix $\mathcal{W}$.
If $n\asymp p$ then the sample covariance matrix $\hat{\Sigma}$ is often-times ill-conditioned, and is furthermore singular when $n<p$. The estimation of either the precision matrix $\Sigma^{-1}$ or the whitening matrix $\mathcal{W}$ is thus challenging in these regimes, especially since $\Sigma^{-1}$ and $\mathcal{W}$ both contains $O(p^{2})$ entries. An universal approach to address this difficulty is to reduce the number of parameters needed for estimating $\Sigma$, for example by assuming that $\Sigma$ has a parametric form with $o(p^{2})$ parameters, or by introducing regularization terms to induce certain structures in the estimate $\hat{\Sigma}$.
The use of regularized estimators for covariance matrices are popular in high-dimensional classification, see e.g., Bickel & Levina (2009); Shao et al. (2011), and theoretical analysis for these estimators are usually based on assumptions about the sparsity of either $\Sigma$ or $\Sigma^{-1}$ and then bounding the estimation error in terms of the spectral norm differences $\|\hat{\Sigma}^{-1}-\Sigma^{-1}\|$ or $\|\hat{\Sigma}-\Sigma\|$. There are, however, two potential drawbacks to this approach. Firstly, quantifying the estimation error in terms of the spectral norm can be quite loose and the resulting bounds might not capture the difference in geometry between the eigenspaces of $\Sigma$ and that of the perturbed matrix $\hat{\Sigma}$ (e.g., Johnstone (2001) observed that the first few largest eigenvalues of the sample covariance matrix $\hat{\Sigma}$ are always larger than those for $\Sigma$ in the high-dimensional setting). Secondly, while it is not always appropriate to assume sparsity of $\Sigma$ or $\Sigma^{-1}$, Cai & Liu (2011) noted that some form of sparsity is needed to guarantee consistent estimation of $\Sigma$ or $\Sigma^{-1}$ under spectral norm error.
In this paper we assume a different structure for $\Sigma$, namely that $\Sigma$ is a spiked-covariance matrix with a few leading eigenvalues (the spike) that are well-separated from the remaining eigenvalues (the bulk). More specifically, we assume that $\Sigma$ satisfies the following condition.
Assumption 1 (Spiked covariance matrix).
Let $\mathbf{u}_{1},\dots,\mathbf{u}_{d}$ be orthonormal vectors in $\mathbb{R}^{p}$ and assume that the covariance matrix $\Sigma$ for the $p$-variate distributions $\mathcal{N}_{p}(\mu_{1},\Sigma)$ and $\mathcal{N}_{p}(\mu_{2},\Sigma)$ is of the form
$\Sigma$ $=\>\sum_{k=1}^{d}\lambda_{k}\mathbf{u}_{k}\mathbf{u}_{k}^{\top}+\sigma^{2}\mathcal{I}_{p}=\mathcal{U}\Lambda\mathcal{U}^{\top}+\sigma^{2}\mathcal{I}_{p}.$ (2.6)
Here $\Lambda=\mathrm{diag}(\lambda_{k})$ is a $d\times d$ diagonal matrix, $\mathcal{U}=(\,\mathbf{u}_{k}\,),\>k\in[d]$ is a $p\times d$ matrix with orthonormal columns, and $\mathcal{I}_{p}$ is the identity matrix. We assume implicitly that $\lambda_{1}\geq\cdots\geq\lambda_{d}>0$, $\sigma>0$ and $d\ll p$.
Covariance matrices with spiked structures have been studied extensively in the high-dimensional statistics literature, see e.g., Johnstone (2001) and Chapter 11 of Yao et al. (2015), and there is a significant number of results for consistent estimation of $\mathcal{U}$ under the assumption that the support of $\mathcal{U}$ is sparse, e.g., either that the number of non-zero rows of $\mathcal{U}$ is small compared to $p$ or that the $\ell_{q}$ quasi-norm, for some $q\in[0,1]$, of the columns of $\mathcal{U}$ are bounded. The case when $q=0$ and $q>0$ correspond to “hard” and “soft” sparsity constraints, respectively; see for example Birnbaum et al. (2012); Berthet & Rigollet (2012); Vu & Lei (2012); Cai et al. (2013) and the references therein. In contrasts to the above cited results, in this paper we do not impose sparsity conditions on $\mathcal{U}$ but instead assume that $\mathcal{U}$ have bounded coherence, namely that the maximum $\ell_{2}$ norm of the rows of $\mathcal{U}$ are of order $O(p^{-1/2})$; see ˜4 for a precise statement. The resulting matrix $\Sigma$ will no longer be sparse. The main rationale for assuming bounded coherence is that the spiked eigenvalues $\Lambda$ can grow linearly with $p$ while still guaranteeing bounded variance in $\Sigma$. There is thus a large gap between the spiked eigenvalues and the bulk eigenvalues, and this justifies the use of PCA as a pre-processing step.
If $\Sigma$ satisfies Condition 1 then the whitening matrix is given by
$\mathcal{W}=\mathcal{U}\mathcal{D}\mathcal{U}^{\top}+\frac{1}{\sigma}(\mathcal{I}_{p}-\mathcal{U}\mathcal{U}^{\top})$ (2.7)
where $\mathcal{D}=\mathrm{diag}(\eta_{k})$ is a $d\times d$ diagonal matrix with diagonal entries $\eta_{k}=(\lambda_{k}+\sigma^{2})^{-1/2}$. The above form for $\mathcal{W}$ suggests the following estimation procedure.
  1. 1.
    Extract the $d$ largest eigenvalues and corresponding eigenvectors of the pooled sample covariance matrix $\hat{\Sigma}$. Let $\hat{\Lambda}$ denote the diagonal matrix of these $d$ largest eigenvalues and $\hat{\mathcal{U}}$ denote the $p\times d$ orthogonal matrix whose columns are the corresponding eigenvectors.
  2. 2.
    Estimate the non-spiked eigenvalues by
    $\hat{\sigma}^{2}=\frac{\mathrm{tr}(\hat{\Sigma})-\mathrm{tr}(\hat{\Lambda})}{p-d}.$ (2.8)
  3. 3.
    Let $\hat{\mathcal{D}}=(\hat{\Lambda}+\hat{\sigma}^{2}\mathcal{I}_{d})^{-1/2}$ and estimate $\mathcal{W}$ by $\hat{\mathcal{W}}=\hat{\mathcal{U}}\hat{\mathcal{D}}\hat{\mathcal{U}}^{\top}+\hat{\sigma}^{-1}(\mathcal{I}_{p}-\hat{\mathcal{U}}\hat{\mathcal{U}}^{\top})$.
Although the sample covariance matrix $\hat{\Sigma}$ is generally a poor estimate of $\Sigma$ when $n\ll p$, the eigenvectors $\hat{\mathcal{U}}$ corresponding to the $d$ largest eigenvalues of $\hat{\Sigma}$ are nevertheless accurate estimates of $\mathcal{U}$. In particular, Fan et al. (2018) and Cape et al. (2019) provide uniform error bound for $\min_{T}\|\hat{\mathcal{U}}T-\mathcal{U}\|_{2\to\infty}$ where the minimum is taken over all $d\times d$ orthogonal transformations $T$; see also Theorem 1 of the current paper. We can thus transform $\hat{\mathcal{U}}$ by an orthogonal transformation $T$ so that the resulting (transformed) rows of $\hat{\mathcal{U}}$ are uniformly close to the corresponding rows of $\mathcal{U}$.
Given the above estimate for $\mathcal{W}$, we then have the prototypical $\mathrm{lda}\circ\mathrm{pca}$ classifier in Algorithm 1. Algorithm 1 contains two tuning parameters, namely $d$, the number of principal components in the PCA step and $s$, the number of coordinates of estimated whitened direction $\hat{\zeta}$ that we preserve. The choices for $d$ and $s$ correspond to the number of spiked eigenvalues in $\Sigma$ and the sparsity level of the whitened direction $\zeta$. We show in Section 3 that, under certain mild conditions, one can consistently estimate $s$ using LABEL:hardthresholding. Consistent estimation of $d$ can be obtained using results in Bai & Ng (2002); Alessi et al. (2010); Hallin & Liška (2007) among others. Algorithm 1 thus yields a classifier that is adaptive with respect to both $d$ and $s$.
Input: $\bar{X}_{1}$, $\bar{X}_{2}$, $\hat{\Sigma}$ and the test sample $\mathbf{Z}$
Output: $\hat{\Upsilon}_{\mathrm{lda}\circ\mathrm{pca}}(\mathbf{Z})$
Algorithm
 // Step 1: Perform PCA on the feature vectors for the training data (standard PCA approach)
 // Step 2: Extract the $d$ largest principal components and obtain $\hat{\mathcal{W}}=\hat{\mathcal{U}}\hat{\mathcal{D}}\hat{\mathcal{U}}^{\top}+\hat{\sigma}^{-1}(\mathcal{I}_{p}-\hat{\mathcal{U}}\hat{\mathcal{U}}^{\top})$.
 // Step 3: Take $\tilde{X}_{i}=\hat{\mathcal{W}}\bar{X}_{i},\>i=1,2$ and $\tilde{X}_{a}=0.5(\tilde{X}_{1}+\tilde{X}_{2})$. Let $\hat{\zeta}=\tilde{X}_{2}-\tilde{X}_{1}$ and form the indices set $\hat{\mathcal{S}}$ by selecting the $s$ largest coordinates of $\hat{\zeta}$ in modulus.
 // Step 4: Given the test sample $\mathbf{Z}$, take $\tilde{\mathbf{Z}}=\hat{\mathcal{W}}\mathbf{Z}$ and plug the sub-vector of $\hat{\zeta}$, $\tilde{\mathbf{Z}}$ and $\tilde{X}_{a}$ corresponding to the indices in $\hat{\mathcal{S}}$ into the Fisher discriminant rule, i.e.,
[…]
Algorithm 1 $\mathrm{lda}\circ\mathrm{pca}$ decision rule
2.4 Related works
2.4.1 Whitening matrix and spiked covariance
We first discuss the relationship between Algorithm 1 and two other relevant classifiers, namely the features annealed independence rule (FAIR) of Fan & Fan (2008) and the LDA with CAT scores (CAT-LDA) of Zuber & Strimmer (2009). FAIR is the independence rule applied to the features prescreened by the two-sample $t$ test while CAT-LDA scores decorrelate the $t$ statistics by first whitening the data using the sample correlation matrix $\mathrm{diag}(\hat{\Sigma})^{-1/2}\,\hat{\Sigma}\,\mathrm{diag}(\hat{\Sigma})^{-1/2}$; CAT-LDA is motivated by the empirical observation that accounting for correlations is essential in the analysis of proteomic and metabolic data.
If there are no correlation or if the correlations are negligible, i.e., when $\Sigma$ is diagonal or approximately diagonal, then Algorithm 1 is essentially equivalent to both FAIR and CAT-LDA. However, when the correlations are not negligible, the performance of IR classifiers such as FAIR degrades significantly due to mis-specification of the covariance structure (Fan et al., 2012). Both Algorithm 1 and CAT-LDA apply a whitening transformation before doing LDA but the choice of whitening matrices are different between the two procedures. In particular CAT-LDA uses the asymmetric whitening transformation $\mathrm{Corr}{(\Sigma)}^{-1/2}{\mbox{diag}}(\Sigma)^{-1/2}$ where $\mathrm{Corr}{(\Sigma)}$ is the matrix of correlations. The choice of whitening transformation is not important if $\Sigma$ is known or if we are in the low-dimensional setting where $p\ll n$ for which consistent estimation of $\Sigma$ is straightforward. It is only when $p\asymp n$ or $p\gg n$ when the different regularization strategies used in CAT-LDA and Algorithm 1 lead to possibly different behaviors for the resulting classifiers; see for example the numerical comparisons given in Section 4 of the current paper. Finally we view the theoretical results in Section 3 as not only providing justification for Algorithm 1 but also serve as examples of theoretical analysis that can be extended to other classifiers in which dimension reduction is done prior to performing LDA. Indeed, to the best of our knowledge, there are no theoretical guarantees for the error rate for the CAT-LDA classifier in the high-dimensional setting.
Finally we note that our paper is not the first work to study high-dimensional classification under spiked covariance matrices of the form in Condition 1. In particular Sifaou et al. (2020a) also analyzed the performance of LDA under spiked covariance structures; however, their theoretical results are premised under different settings and framework from ours and most importantly they do not show that their classification rule achieves the Bayes error rate; see Remark 4 for detailed comparisons of the assumptions and theoretical results in our paper against that of Sifaou et al. (2020a). With a different focus, Hao et al. (2015) examined the effects of spiked eigenvalues when trying to find an orthogonal transformation $T$ of the discriminant direction $\beta$ so that $T\beta$ is sparse. More specifically they show that if $\Sigma$ is known then $T$ can be computed using the eigen-decomposition of the matrix $\Sigma_{\mathrm{rot}}$ where
$\Sigma_{\mathrm{rot}}=\Sigma+\gamma\Delta\mu\Delta\mu^{\top},\qquad\text{for a given $\gamma>0$}$ (2.10)
where $\Delta\mu=\mu_{2}-\mu_{1}$ is the vector of mean differences between the two classes. Let $\mathcal{U}_{\mathrm{rot},m}$ be the $p\times m$ matrix whose columns consist of the orthonormal eigenvectors of $\Sigma_{\mathrm{rot}}$ corresponding to the $m$ largest eigenvalues. When $m=p$ then $\mathcal{U}_{\mathrm{rot},p}$ diagonalizes $\Sigma_{\mathrm{rot}}$ and Hao et al. (2015) showed that $\mathcal{U}_{\mathrm{rot},p}$ further sparsifies $\beta$ in that $\|\mathcal{U}_{\mathrm{rot},p}^{\top}\,\beta\|_{0}\leq d+1$ whenever $\Sigma$ satisfies the spiked covariance assumption in ˜1, and hence it might be beneficial to rotate the data before performing classification.
When $\Sigma$ is unknown Hao et al. (2015) propose the following procedure for estimating $\mathcal{U}_{\mathrm{rot},m}$ for some choice of $m\leq\min\{n,p\}$; here $\gamma>0$ is a user-specified parameter.
Algorithm
 // Step 1: Perform PCA on $\hat{\Sigma}_{\mathrm{rot}}=\hat{\Sigma}+\gamma(\bar{X}_{2}-\bar{X}_{1})(\bar{X}_{2}-\bar{X}_{1})^{\top}$ to extract the $m$ largest principal components. Let $\hat{\mathcal{U}}_{\mathrm{rot},m}$ be the resulting $p\times m$ matrix.
 // Step 2: Rotate the training set to $\{\hat{\mathcal{U}}_{\mathrm{rot},m}^{\top}X_{i1},\dots,\hat{\mathcal{U}}_{rot,m}^{\top}X_{in_{i}}\}$ where $\hat{\mathcal{U}}_{\mathrm{rot},m}^{\top}X_{ij}\in\mathbb{R}^{m}$ for $i\in\{1,2\}$ and $j\in\{1,2,\dots,n_{i}\}$.
 // Step 3: Apply some discriminant direction based LDA method such as Fan et al. (2012); Cai & Liu (2011); Mai et al. (2015) on the rotated data set $\{\hat{\mathcal{U}}_{\mathrm{rot},m}^{\top}X_{ij}\}$.
Algorithm 2 Rotation as a preprocessing in LDA
We will use this pre-processing step in some of our numerical experiments in Section 4. Nevertheless we emphasize that the theoretical results of Hao et al. (2015) assume $\Sigma$ is known as they are mainly concerned with the analysis of different approaches for sparsifying $\beta$, i.e., they explicitly chose not to address the important issue of how the estimation error for $\hat{\mathcal{U}}_{\mathrm{rot},m}$ impacts the classification accuracy.
2.4.2 Connection to high-dimensional sparse LDA
Let $\mathbf{X}$ be the $(n_{1}+n_{2})\times p$ matrix whose rows are the $\{X_{ij}\}$ and $\mathcal{Y}\in\mathbb{R}^{n}$ be the vector whose first $n_{1}$ elements are set to $(n/n_{1})$ and the remaining $n_{2}$ elements are set to $n/n_{2}$. Mai et al. (2012) reframed LDA in high-dimension as the solution to a Lasso-type problem (we have omitted the intercept term for simplicity of presentation)
$\hat{\beta}$ $=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\;\frac{1}{2}\|\mathcal{Y}-\mathbf{X}\beta\|^{2}+\lambda\|\beta\|_{1}$ (2.11)
for some $\lambda>0$. In a similar spirit to Eq.˜2.11, we might, conceptually, also reformulate Algorithm 1 as a (general) Lasso-type problem (Tibshirani & Taylor (2011))
$\hat{\beta}$ $=\operatorname*{arg\,min}_{\beta\in\mathbb{R}^{p}}\;\frac{1}{2}\|\mathcal{Y}-\mathbf{X}\beta\|^{2}+\lambda\|\Psi\beta\|_{1}$ (2.12)
for some $\lambda>0$, where $\Psi=\Sigma^{1/2}$ (once again omitting the intercept for simplicity). As $\Sigma$ is invertible, Eq.˜2.12 is equivalent to first solving
$\hat{\zeta}$ $=\operatorname*{arg\,min}_{\zeta\in\mathbb{R}^{p}}\;\frac{1}{2}\|\mathcal{Y}-\tilde{\mathbf{X}}\zeta\|^{2}+\lambda\|\zeta\|_{1}$ (2.13)
where $\tilde{\mathbf{X}}=\mathbf{X}\Sigma^{-1/2}$, and then setting $\hat{\beta}=\Sigma^{-1/2}\hat{\zeta}$. Comparing Eqs.˜2.12 and 2.13 against Eq.˜2.11, one could argue that the main difference between $\mathrm{lda}\circ\mathrm{pca}$ in Algorithm 1 and the lassoed LDA classifier of Mai et al. (2012) is due to the different choice of the predictor variables $\mathbf{X}$ vs $\tilde{\mathbf{X}}$ in the Lasso regression. This argument, however, overlooks the important fact that the transformation $\Sigma^{-1/2}$ needs to be estimated in Algorithm 1 while for the general Lassso in Eq.˜2.12, the matrix $\Psi$ is specified a priori and thus free from any estimation error. Indeed, while there have been significant efforts devoted to understanding the solution path of Eq.˜2.12 in the high-dimensional setting – see, for example, Duan et al. (2016); Tibshirani & Taylor (2011) among others – its behavior when $\Psi$ is determined empirically from the data is yet to be theoretically investigated.
Next we note that, instead of Eq.˜2.11, one can also estimate $\zeta$ via an $\ell_{1}$ optimization problem similar to that of the Dantzig selector (Candes & Tao, 2005; Cai & Liu, 2011; Cai & Zhang, 2019), namely
$\hat{\zeta}=\operatorname*{arg\,min}_{\zeta\in\mathbb{R}^{p}}\Big\{\|\zeta\|_{1}\,\,\text{subject to }\>\|\hat{\Sigma}^{1/2}\zeta-(\bar{X}_{2}-\bar{X}_{1})\|_{\infty}\leq\lambda\Big\}.$ (2.14)
The main difference between Algorithm 1 and Eq.˜2.14 is that Algorithm 1 first whitens the feature vectors using an empirical estimate of the covariance matrix followed by feature selection on the whitened data (see Eq.˜2.13). In contrast, Eq.˜2.14 attempts to find a $\zeta$ for which its transformation $\hat{\Sigma}^{1/2}\zeta$ is most similar to $\bar{X}_{2}-\bar{X}_{1}$. While the approaches underlying Algorithm 1 and Eq.˜2.14 are quite similar, we believe that Algorithm 1 provides a more direct link between PCA and LDA for high-dimensional classification as it is particularly suitable for high-dimensional data generated from a (possibly low-dimenisonal) factor model (where the covariance matrix will now have a spiked structure with leading eigenvalues that grow with the dimension $p$). See Fan et al. (2021) for further discussion of factor models and their applications to statistics and machine learning.
Although Algorithm 1 and Eq.˜2.11 share a conceptual link in optimization formulation, they differ significantly in algorithmic structure, leading to distinct computational profiles. We now compare the computational complexity of $\mathrm{lda}\circ\mathrm{pca}$ and Lassoed LDA. For our method, the dominant cost comes from computing the top $\bar{d}$ singular values of the centered data matrix $\mathbf{X}$, with cost $O(np\bar{d})$, where $\bar{d}$ upper-bounds the rank $d$; see Halko et al. (2011) and Feng & Yu (2023) for more details. In contrast, Lassoed LDA solves an $\ell_{1}$-penalized least squares problem with complexity $O(npT)$, where $T$ is the number of iterations until convergence. The iteration count $T$ can be substantial in practice, especially when regularization is weak or $\mathbf{X}$ is ill-conditioned as often occurs in settings with low-rank or highly correlated features. Consequently, our method provides substantial computational advantages in high-dimensional latent variable models, while maintaining competitive classification accuracy and working in a reduced-dimensional space (see Section˜4).
Finally we conclude this section by comparing Algorithm 1 with the principal component classifiers proposed in Bing et al. (2024); Bing & Wegkamp (2023). In particular both Bing & Wegkamp (2023) and Bing et al. (2024) assumed that the discriminant direction $\beta$ (and the whitening direction $\zeta$ as in our work) lies entirely within the column space spanned by the $d$ leading principal components $\mathcal{U}$. However, as noted by Jolliffe (1982), the principal components associated with small eigenvalues can be just as important as those associated with large eigenvalues in real data analysis (see also the simulation settings for Model 1 and Model 3 in Section˜4). By working with the whitened data, Algorithm 1 also takes into account the non-leading principal components and furthermore, by performing feature selection in the whitened space, avoid the need to specify a priori the low-dimensional subspaces containing $\beta$ and/or $\zeta$.
3 Theoretical properties
In this section, we derive the theoretical properties of the $\mathrm{lda}\circ\mathrm{pca}$ classifier in Algorithm 1 for the case where the feature vectors $X$ are sampled from a mixture of two multivariate Gaussians. Extensions of these results to the case where $X$ is a mixture of $K\geq 2$ elliptical, but not necessarily multivariate normal, distributions are discussed in Section 7.3 and Section 5.1.
If $X\sim\pi_{1}\mathcal{N}_{p}(\mu_{1},\Sigma)+(1-\pi_{1})\mathcal{N}_{p}(\mu_{2},\Sigma)$ then the Fisher’s rule $\Upsilon_{F}$ has error rate
$R_{F}=\pi_{1}\Phi\Bigl(-\frac{1}{2}\|\zeta\|+\|\zeta\|^{-1}\ln\frac{1-\pi_{1}}{\pi_{1}}\Bigr)+(1-\pi_{1})\Phi\Bigl(-\frac{1}{2}\|\zeta\|-\|\zeta\|^{-1}\ln\frac{1-\pi_{1}}{\pi_{1}}\Bigr).$ (3.1)
Here $\Phi$ is the cumulative distribution function for $\mathcal{N}(0,1)$ and $\zeta=\Sigma^{-1/2}(\mu_{2}-\mu_{1})$; see e.g., Ripley (1996, Section 2.1). We note that $R_{F}$ is the smallest mis-classification error achievable by any classifier and, furthermore, is monotone decreasing as $\|\zeta\|$ increases. If $\left\lVert\zeta\right\rVert\rightarrow 0$ then $R_{F}\rightarrow\min\{\pi_{1},1-\pi_{1}\}$ and $\Upsilon_{F}$ is no better than assigning every data point to the most prevalent class, while if $\left\lVert\zeta\right\rVert\rightarrow\infty$ then $R_{F}\rightarrow 0$ and $\Upsilon_{F}$ achieves perfect accuracy. The cases where $R_{F}\rightarrow 0$ or $R_{F}\rightarrow\min\{\pi_{1},1-\pi_{1}\}$ are, theoretically, uninteresting and hence in this paper we only focus on the case where $0<\|\zeta\|<\infty$. We therefore make the following assumption.
Assumption 2.
Let $S_{\zeta}$ denote the set of indices $i$ for which $\zeta_{i}\not=0$. Also let $\mathcal{C}_{0}>0$, $M>0$ and $\mathcal{C}_{\zeta}>0$ be constants not depending on $p$ such that $s_{0}:=|\mathcal{S}_{\zeta}|\leq M$ and
$\min_{j\in\mathcal{S}_{\zeta}}|\zeta_{j}|\geq\mathcal{C}_{0},\qquad\max\bigl\{\|\Sigma^{-1/2}\mu_{1}\|,\|\Sigma^{-1/2}\mu_{2}\|\bigr\}\leq\mathcal{C}_{\zeta}.$
In addition, we assume that the number of spikes $d$ in the spiked covariance model (see ˜1) is fixed and does not grow with $p$ or $n$.
Remark 1.
˜2 implies $0<\mathcal{C}_{0}\sqrt{s_{0}}\leq\|\zeta\|\leq 2\mathcal{C}_{\zeta}<\infty$. We emphasize that sparsity is imposed on the whitened direction so that only a few transformed features contribute to classification outcome; similar assumptions can be found in Silin & Fan (2022). To further explain this condition, we note that sparsity of the discriminant direction $\beta=\Sigma^{-1}(\mu_{1}-\mu_{2})$ also implies that only a small subset of the raw covariates affects the response (as the classification boundary for a feature vector $\bm{x}$ is given by $\bm{1}\{\bm{x}^{\top}\beta>a\}$ for some $a\in\mathbb{R}$). However, as noted by Zhu & Bradic (2018) and Hall et al. (2014), sparsity of $\beta$ is incompatible with many real data application such as genome-wide gene expression profiling where all genes are believed to play a role in disease markers, or analysis of micro-array data to identify leukemia or colon/prostate cancer. In contrast sparsity of the whitened direction $\zeta=\Sigma^{-1/2}(\mu_{1}-\mu_{2})$ still allows for $\beta=\Sigma^{-1/2}\zeta$ to be non-sparse, thereby circumventing the issue described above. Lastly, throughout our theoretical results and proofs, we assume that the sparsity level $s_{0}$ is fixed and independent of the sample size $n$, though it may be arbitrary. The case where $s_{0}$ grows with $n$ can be handled mutatis mutandis. However, it requires more careful control of technical arguments and is therefore left to the interested reader.
Our theoretical results are large-sample results in which the sample sizes $n_{1}$ and $n_{2}$ for the training data increase as the dimension $p$ increases, and thus our next assumption specifies the asymptotic relationships between these quantities and the eigenvalues of $\Sigma$.
Assumption 3.
Let $\sigma>0$ be fixed and suppose that
$\frac{n_{1}}{n_{2}}=\Theta(1),\quad\ln{p}=o(n).$
Furthermore, the spiked eigenvalues $\lambda_{1},\dots,\lambda_{d}$ of $\Sigma$ satisfy
$\lambda_{k}=\Theta(p),\qquad\text{for all $k\in[d]$}.$
If $\Sigma$ satisfies ˜3 then $\mathbf{tr}(\Sigma)=\Theta(p)$ and $\bm{r}(\Sigma)=\mathrm{tr}(\Sigma)/\|\Sigma\|=\mathcal{O}(1)$. Recall that $\bm{r}(\Sigma)$ is the effective rank of $\Sigma$. ˜3 also implies that the spiked eigenvalues of $\Sigma$ are unbounded as $p$ increases and this assumption distinguishes our theoretical results from existing results in the literature wherein it is generally assumed that the eigenvalues of $\Sigma$ are bounded; see Bickel & Levina (2004); Fan & Fan (2008); Shao et al. (2011); Cai & Zhang (2019) for a few examples of results under the bounded eigenvalues assumption. While the bounded eigenvalues assumption is prevalent, it can also be problematic for high-dimensional data as it ignores the strong signals present in many real data applications; see Fan et al. (2013) and Wang & Fan (2017) for further discussions of this issue. Indeed, a standard heuristic for PCA is to keep the $d$ largest principal components that explains $90\%$ or $95\%$ of the variability in the data, and hence if $p$ is large and $\Sigma$ has bounded eigenvalues then one has to choose $d=\Theta(p)$ which is inconsistent with the use of PCA as a dimension reduction procedure.
If $\Sigma$ has a spiked eigenvalue structure as in ˜1 then $\mathcal{W}$ is given by
$\mathcal{W}=\Sigma^{-1/2}=\mathcal{U}(\Lambda+\sigma^{2}\mathcal{I})^{-1/2}\mathcal{U}^{\top}+\sigma^{-1}(\mathcal{I}-\mathcal{U}\mathcal{U})^{\top}$ (3.2)
and a natural estimate for $\mathcal{W}$ is
$\hat{\mathcal{W}}=\hat{\mathcal{U}}(\hat{\Lambda}+\hat{\sigma}^{2}\mathcal{I})^{-1/2}\hat{\mathcal{U}}^{\top}+\hat{\sigma}^{-1}(\mathcal{I}-\hat{\mathcal{U}}\hat{\mathcal{U}}^{\top})$ (3.3)
where $\hat{\Lambda}$ is the diagonal matrix containing the $d$ largest eigenvalues of the pooled covariance matrix, $\hat{\mathcal{U}}$ is the $p\times d$ orthonormal matrix whose columns are the corresponding eigenvectors, and $\hat{\sigma}^{2}$ is as defined in Eq.˜2.8. We now make the following assumption on $\mathcal{U}$.
Assumption 4 (Bounded Coherence).
There is a constant $\mathcal{C}_{\mathcal{U}}\geq 1$ independent of $n$ and $p$ such that
$\left\lVert\mathcal{U}\right\rVert_{2\rightarrow\infty}\leq\frac{\mathcal{C}_{\mathcal{U}}\sqrt{d}}{\sqrt{p}}.$
Remark 2.
The bounded coherence assumption appears frequently in statistical inference for matrix-valued data. More specifically, as every column in $\mathcal{U}$ has $\ell_{2}$ norm equal to $1$, the rows $\mathcal{U}$ have, on average, $\ell_{2}$ norm of order $\sqrt{d/p}$. ˜4 then guarantees that the maximum $\ell_{2}$ norm of the rows of $\mathcal{U}$ is also of order $\sqrt{d/p}$. For more discussions about the bounded coherence assumption in the context of matrix completion, covariance matrix estimation, and random matrix theory, see Candès & Recht (2009); Fan et al. (2018); Rudelson & Vershynin (2016); Bloemendal et al. (2014) among others. Finally we note that if $\mathcal{U}$ satisfies the bounded coherence assumption with constant $\mathcal{C}_{\mathcal{U}}$ then
$\Sigma_{ii}\leq\,\frac{\mathcal{C}_{\mathcal{U}}^{2}\lambda_{1}d}{p}+\sigma^{2},\qquad\text{for all $i\in[p]$},$
which together with ˜3 implies $\max_{i\in[p]}\Sigma_{ii}=\mathcal{O}(1).$ In summary, ˜4 allows for the spiked eigenvalues $\lambda_{1},\dots,\lambda_{d}$ of the covariance matrix $\Sigma$ to grow linearly with $p$ while also guaranteeing that the entries of $\Sigma$ remains bounded, i.e., each variable in $X_{ij}$ has a finite variance.
We next state a result on the estimation accuracy of $\hat{\mathcal{U}}$. This result is a slight extension of an earlier result in Cape et al. (2019). More specifically, Cape et al. (2019) assume $\mathbb{E}[\mathbf{X}]=0$ and hence the sample covariance matrix is simply $\tfrac{1}{n}\mathbf{X}^{\top}\mathbf{X}$. In this paper we used the pooled sampled covariance matrix which requires first centering the feature vectors by the sample means of each class.
Theorem 1.
Let $\mathbf{X}$ be a $n\times p$ matrix where the rows $X_{1},\dots,X_{n}$ are i.i.d samples from $\pi_{1}\mathcal{N}(\mu_{1},\Sigma)+(1-\pi_{1})\mathcal{N}(\mu_{2},\Sigma)$ and $\Sigma$ satisfies ˜1, ˜3, and ˜4. Let $\hat{\mathcal{U}}$ be the matrix of eigenvectors corresponding to the $d$ largest eigenvalues of the pooled sample covariance matrix $\hat{\Sigma}$. Then there exists a $d\times d$ orthogonal matrix $\Xi_{\mathcal{U}}$ and a constant $C>0$ such that with probability at least $1-\mathcal{O}(p^{-2})$,
$\|\hat{\mathcal{U}}-\mathcal{U}\Xi_{\mathcal{U}}\|_{2\rightarrow\infty}\leq\mathcal{C}\sqrt{\frac{d^{3}\ln{p}}{np}}.$ (3.4)
Note that, for simplicity, we assume in ˜1 as well in the subsequent part of this paper that $d$, the number of spiked eigenvalues, is known. If $d$ is unknown then it can be consistently estimated using the ratio between consecutive eigenvalues, similar to the procedures in Ahn & Horenstein (2013). More specifically, from ˜3 and ˜1 we have with high probabliity that $\hat{\lambda}_{k}=\Theta(p)$ for $k\leq d$ and $\hat{\lambda}_{k}=O(\sigma^{2})$ for $k\geq d+1$. Here $\hat{\lambda}_{k}$ are the eigenvalues of $\hat{\Sigma}$. There thus exists a significant gap between $\hat{\lambda}_{d-1}/\hat{\lambda}_{d}=O(1)$ and $\hat{\lambda}_{d}/\hat{\lambda}_{d+1}=\Omega(p)$, and hence, letting $\hat{d}$ be the smallest index $k$ for which $\hat{\lambda}_{k}/\hat{\lambda}_{k+1}=\Omega(\ln p)$, we have $\hat{d}=d$ asymptotically almost surely.
We now analyze the classification accuracy of $\mathrm{lda}\circ\mathrm{pca}$. Recall that the main idea behind $\mathrm{lda}\circ\mathrm{pca}$ is that we first construct an estimate $\hat{\zeta}$ for the whitened direction $\zeta$ and then project our whitened data onto the $d$ largest coordinates, in magnitude, of $\hat{\zeta}$ (see Algorithm 1). By Eq.˜3.1, the Bayes error rate $R_{F}$ is a monotone decreasing function of $\|\zeta\|$ and hence, to achieve $R_{F}$ it is only necessary to recover the indices in $S_{\zeta}=\{i\colon\zeta_{i}\not=0\}$. In summary the error rate for $\mathrm{lda}\circ\mathrm{pca}$ converges to $R_{F}$ provided we can (1) bound the error $\hat{\zeta}-\zeta$ and (2) show that thresholding $\hat{\zeta}$ perfectly recovers $S_{\zeta}$.
Theorem 2.
Under Assumptions1-4, there exists a constant $C>0$ such that with probability at least $1-\mathcal{O}(p^{-2})$,
$\|\hat{\zeta}-\zeta\|_{\infty}\leq C\sqrt{\frac{\ln{p}}{n}}.$ (3.5)
˜2 implies $\zeta_{i}\not=0$ if and only if $\zeta_{i}>\mathcal{C}_{0}$ for some constant $\mathcal{C}_{0}>0$. By Theorem 2, if $\zeta_{i}=0$ then $\hat{\zeta}_{i}=O((n^{-1}\ln p)^{1/2})$ with high probability. Let $\tilde{\zeta}$ be a hard thresholding of $\hat{\zeta}$, namely
$\tilde{\zeta}_{j}=\hat{\zeta}_{j}\mathbbm{1}(|\hat{\zeta}_{j}|>t_{n}),\qquad j\in[p]$ (3.6)
where $t_{n}=(n^{-1}\ln{p})^{\alpha}$ for some $0<\alpha<\frac{1}{2}$. Given $\tilde{\zeta}$, define the active set $\tilde{S}=\{j:\tilde{\zeta}_{j}\neq 0\}$. The mis-classification rate for $\mathrm{lda}\circ\mathrm{pca}$ conditional on the training data $\{X_{11},\dots,X_{1n_{1}}\}$ and $\{X_{11},\dots,X_{1n_{1}}\}$, is then
$\hat{R}_{\mathrm{lda}\circ\mathrm{pca}}=\mathbb{P}(\text{label}(\mathbf{Z})\neq\hat{\Upsilon}_{\mathrm{lda}\circ\mathrm{pca}}(\mathbf{Z})\mid\{X_{11},\dots,X_{1n_{1}}\},\{X_{21},\dots,X_{2n_{2}}\})$ (3.7)
The following result shows that $\tilde{S}$ recovers $S_{\zeta}$ exactly and $\mathrm{lda}\circ\mathrm{pca}$ is asymptotically Bayes-optimal.
Theorem 3.
Suppose that $\mathbf{Z}\sim\pi_{1}\mathcal{N}_{p}(\mu_{1},\Sigma)+(1-\pi_{1})\mathcal{N}_{p}(\mu_{2},\Sigma)$ where $\pi_{1}\in(0,1)$. Suppose ˜1 through ˜4 are satisfied. We then have
$\mathbb{P}(\tilde{S}\neq\mathcal{S}_{\zeta})=\mathcal{O}(p^{-2}).$ (3.8)
[…]
$|\hat{R}_{\mathrm{lda}\circ\mathrm{pca}}-R_{F}|\longrightarrow 0$ (3.9)
almost surely as $n,p\rightarrow\infty$.
Our theoretical framework and results can be extended to elliptical distributions with minimal modifications; see the discussion in Section 7.3. Key remarks on our main results are presented below.
Remark 3.
Theorem 2 and Theorem 3 appear, at first blush, quite similar to the results in Mai et al. (2012) for lassoed LDA using the discriminant direction $\beta=\Sigma^{-1}(\mu_{1}-\mu_{2})$. However the underlying assumptions behind these results are very different. In particular lassoed LDA relies on irrepresentable condition analogous to those in Zhao & Yu (2006) to achieve consistent variable selection, and thus the discriminative variables cannot be highly correlated with the remaining (irrelevant) variables. In contrast ˜3 do not assume an irrepresentable condition and furthermore ˜1 only assume boundedness of the non-spiked eigenvalues but allows for diverging spiked eigenvalues.
Remark 4.
We now compare our results and settings with those in Sifaou et al. (2020a). Firstly, Sifaou et al. (2020a) assumed: (1) proportional growth rate $p/n\rightarrow c$ for some finite constant $c>0$, (2) distinct spiked eigenvalues $\lambda_{1}>\lambda_{2}>\dots>\lambda_{d}$ and lastly (3) $\|\Sigma\|=\mathcal{O}(1)$. In contrast, our work is derived under a more challenging set-up wherein (1) the dimension $p$ and the number of training samples $n$ satisfy $n^{-1}\ln{p}\rightarrow 0$ (˜3), (2) some or even all of the eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{d}>0$ can be equal (˜1) and (3) the leading eigenvalues $\{\lambda_{k}\}_{k=1}^{d}$ can diverge with $n$ (˜3). We emphasize that $n^{-1}\ln p\rightarrow 0$ allows for $p=n^{\gamma}$ for any $\gamma>1$, so that $p/n\rightarrow\infty$.
The assumptions used in Sifaou et al. (2020a) are because their theoretical analysis rely heavily on existing results in random matrix theory as presented in Donoho & Ghorbani (2018); Donoho et al. (2018), which generally require $p/n<\infty$ as well as bounded eigenvalues. In contrast, if $p/n\rightarrow\infty$, then Donoho & Ghorbani (2018) shows that the best estimate of the true covariance matrix is typically a diagonal matrix (see also Bickel & Levina (2004)), which might make the results in Sifaou et al. (2020a) sub-optimal for the high-dimensional regime $p/n\rightarrow\infty$ considered in our paper.
Finally, Sifaou et al. (2020a) did not guarantee that their classification rule is Bayes-optimal. Rather, they only show that the its mis-classification rate converge to some expression given in Theorem 3 of Sifaou et al. (2020a), and is strictly larger than Bayes error when $p/n$ increases. Sifaou et al. (2020b) subsequently extended the results in Sifaou et al. (2020a) to the case of heterogeneous covariance matrices but the resulting classifier is once again not Bayes-optimal. In contrast, Theorem 3 and Theorem 6 showed that the error rate for $\mathrm{lda}\circ\mathrm{pca}$ and $\mathrm{qda}\circ\mathrm{pca}$ (Quadratic Discriminant Analysis, QDA) converge to the Bayes error rate in the case of equal and unequal (class-conditional) covariance matrices, respectively. These results are, to the best of our knowledge, are among the first to show Bayes consistency when combining PCA with LDA/QDA for high-dimensional classification under divergent spikes. A key reason for this improvement lies in two elements: (1) the assumption of sparsity in the signal vector $\zeta$, which mitigates estimation error high-dimensional means and covariances (motivated by Tibshirani et al. (2002); Fan & Fan (2008)) and (2) the use of the entrywise matrix perturbation bounds, providing precise control over the estimated eigenspace $\hat{\mathcal{U}}$ and the resulting whitened direction $\hat{\zeta}$.
4 Numerical results
We now present simulation results and real data analysis for $\mathrm{lda}\circ\mathrm{pca}$. Comparisons will be made against the nearest shrunken centroids method (NSC) of Tibshirani et al. (2002), sparse linear discriminant analysis (SLDA) of Shao et al. (2011), direct sparse discriminant analysis (DSDA) of Mai et al. (2012), adaptive linear discriminant analysis (AdaLDA) of Cai & Zhang (2019), and LDA rule with CAT scores (CAT-LDA) of Zuber & Strimmer (2009); note that CAT-LDA uses a different sphering transformation compared to $\mathrm{lda}\circ\mathrm{pca}$. There are three other commonly used classifiers that are not included in our comparisons, namely the naive Bayes rule, the linear programming discriminant (LPD) (Cai & Liu, 2011) and the regularized optimal affine discriminant (ROAD)(Fan et al., 2012). We omit these classifiers because (1) the naive Bayes rule is a special case of the NSC rule without soft thresholding of the mean vectors (2) Mai & Zou (2013) showed that the ROAD and DSDA classifiers are equivalent and (3) Cai & Zhang (2019) showed that the AdaLDA rule is a refinement of the LPD rule, i.e., compared to the LPD rule, the AdaLDA rule allows for “heteroscedastic constraints” and requires no tuning parameters.
Implementations of the DSDA and CAT-LDA rules are based on the TULIP and sda library in $\mathtt{R}$ while the implementation of the AdaLDA rule is based on $\mathtt{Matlab}$ codes provided in ADAM github repository. We also consider the rotation pre-processing step of Hao et al. (2015), which yields the transformed data $\{\mathcal{U}_{\mathrm{rot},n}^{\top}X_{11},\dots,\mathcal{U}_{\mathrm{rot},n}^{\top}X_{1n_{1}}\}$ and $\{\mathcal{U}_{\mathrm{rot},n}^{\top}X_{21},\dots,\mathcal{U}_{\mathrm{rot},n}^{\top}X_{2n_{2}}\}$ based on $\mathtt{Matlab}$ codes from HDRotation where we set $\gamma=0.25$ in Eq.˜2.10. These transformed data are then used as input to the DSDA and AdaLDA rules; we denote the resulting classifiers as DSDA(rot) and AdaLDA(rot).
4.1 Simulated examples
We consider three different simulation settings. For each setting, the number of features is set to $p=800$ and we generate $n_{1}=n_{2}=100$ data points from each class for the training data and also generate $n_{1}=n_{2}=100$ data points from each class for the testing data. The classification accuracy of the classifiers in each simulation setting are computed based on $200$ Monte Carlo replications. The mean vectors of the two classes are $\mu_{1}=\mathbf{0}_{800}$ and $\mu_{2}=(\bm{1}_{10},\mathbf{0}_{790})$, i.e., the vector $\mu_{2}$ contains $10$ non-zero entries with values all equal to $1$. We consider the following models for the covariance matrix $\Sigma$. These models were considered previously in Fan et al. (2012); Mai et al. (2012); Cai & Zhang (2019); Cai & Liu (2011), among others. Throughout this sub-section, the number of selected features refers to the dimensionality of the representation on which each method operates. In particular, $\mathrm{lda}\circ\mathrm{pca}$ performs feature selection in the whitened feature space, whereas other methods select features directly from the raw feature space. See Section˜7.2 for a detailed study in which the sparsity levels of the mean difference $\mu_{2}-\mu_{1}$, the discriminant direction $\beta$, and the whitened direction $\zeta$ are matched.
  1. 1.
    Model 1 (equal correlation): Here $\Sigma=(\sigma_{ij})_{p\times p}=\rho\bm{1}\bm{1}^{\top}+(1-\rho)\mathcal{I}_{p}$, i.e., $\sigma_{ij}=\rho$ for $i\neq j$ and $\sigma_{ij}=1$ for $i=j$. With this covariance structure $\Sigma$, the discriminant direction $\beta=\Sigma^{-1}(\mu_{1}-\mu_{0})$ and whitened direction $\zeta=\Sigma^{-1/2}(\mu_{1}-\mu_{0})$ are non-sparse (all entries of $\beta$ and $\zeta$ are non-zero).
  2. 2.
    Model 2 (block diagonal with equal correlation) Here $\Sigma=(\sigma_{ij})_{p\times p}$ is assumed to be a block diagonal matrix with two blocks of size $20\times 20$ and $(p-20)\times(p-20)$. Both diagonal blocks are also of the form $\rho\bm{1}\bm{1}^{\top}+(1-\rho)\mathcal{I}$ where the correlation $\rho$ is the same for both blocks. The discriminant direction $\beta$ and whitened direction $\zeta$ are sparse in this model, with $\beta$ and $\zeta$ both having $20$ non-zero entries.
  3. 3.
    Model 3 (random correlation) Here $\Sigma=\mathcal{L}\mathcal{L}^{\top}+c_{\mathcal{L}}\mathcal{I}_{p}$ where $\mathcal{L}\in\mathbb{R}^{p\times 10}$ with $\mathcal{L}_{ij}$ generated from $\mathcal{N}(0,1)$ and $c_{\mathcal{L}}=\min_{i\in[p]}[\mathcal{L}\mathcal{L}^{\top}]_{ii}$. Note that we generate a new $\mathcal{L}$ for every Monte Carlo replicate. For further comparison, we also consider $\mathcal{L}_{ij}$ generated from the uniform distribution on [1,1][-1,1] and the Student’s t-distribution with $5$ degrees of freedom. The discriminant direction $\beta$ and $\zeta$ are generally non-sparse in this model, i.e., all of their entries are non-zero.
[…]
For data generated according to Model 1, Table 1 and Table 2 show that $\mathrm{lda}\circ\mathrm{pca}$ achieves the highest accuracy while using only a small number of features compared to the other classifiers. Note that although CAT-LDA also apply a whitening transformation before performing LDA, the accuracy of CAT-LDA is much worse compared to that of $\mathrm{lda}\circ\mathrm{pca}$. The DSDA and DSDA(rot) classifiers have slightly better accuracies compared to those for the AdaLDA and AdaLDA(rot) classifiers; recall that the DSDA(rot) and AdaLDA(rot) classifiers first applied the rotation pre-processing step of Hao et al. (2015) before running DSDA and AdaLDA on the transformed data. The NSC classifier has the largest mis-classification error; this is a consequence of the NSC rule ignoring the correlation structure in $\Sigma$. Finally, the oracle classifier correspond to the LDA rule where $\pi_{1},\mu_{0},\mu_{1}$ and $\Sigma$ are known, and thus its error rate is the Bayes error rate from Eq.˜3.1.
Table 1: Mis-classification rate ($\%$) with standard deviations ($\%$) in parentheses for equal correlation setting (model 1), based on $200$ independent Monte Carlo replicates.
$\rho$ 0.50 0.60 0.70 0.80 0.90
Oracle 1.37 (0.87) 0.61 (0.53) 0.19 (0.30) 0.19 (0.30) 0.00 (0.00)
$\mathrm{lda}\circ\mathrm{pca}$ 1.74 (1.00) 1.00 (0.82) 0.55 (0.67) 0.55 (0.67) 0.22 (0.39)
CAT-LDA 7.18 (2.22) 4.70 (1.77) 2.60 (1.34) 1.10 (0.88) 0.38 (0.56)
DSDA 3.27 (1.47) 1.96 (1.08) 0.74 (0.62) 0.76 (0.65) 0.00 (0.04)
DSDA(rot) 6.31 (1.83) 3.26 (1.27) 1.16 (0.83) 0.30 (0.41) 0.01 (0.06)
AdaLDA 4.15 (1.61) 2.62 (1.24) 1.22 (0.91) 0.31 (0.42) 0.00 (0.04)
AdaLDA(rot) 7.04 (2.07) 3.70 (1.26) 1.19 (0.79) 0.13 (0.26) 0.00 (0.00)
SLDA 17.97 (3.53) 14.46 (2.99) 11.14 (2.43) 11.14 (2.43) 1.64 (0.92)
NSC 20.38 (8.53) 22.60 (8.38) 24.64 (8.36) 24.64 (8.36) 29.26 (8.04)
Table 2: Average number of nonzero coefficients with standard deviations in parentheses for equal correlation setting (model 1), based on $200$ independent Monte Carlo replicates.
$\rho$ 0.50 0.60 0.70 0.80 0.90
$\mathrm{lda}\circ\mathrm{pca}$ 12.04 (4.53) 11.31 (4.10) 9.52 (4.01) 9.52 (4.01) 3.68 (0.98)
CAT-LDA 24.85 (3.74) 24.70 (3.58) 24.60 (3.97) 22.29 (3.93) 15.45 (2.39)
DSDA 96.20 (31.06) 106.42 (32.57) 117.15 (30.52) 117.05 (29.97) 96.48 (8.66)
DSDA(rot) 33.78 (32.71) 36.88 (38.13) 57.23 (61.02) 147.57 (66.97) 176.29 (5.08)
AdaLDA 46.54 (5.95) 45.84 (5.25) 46.72 (6.41) 47.86 (5.88) 48.65 (5.02)
AdaLDA(rot) 5.68 (1.82) 5.57 (1.84) 5.86 (1.82) 7.83 (2.54) 19.85 (9.96)
SLDA 728.84 (226.63) 788.16 (95.69) 799.96 (0.21) 799.96 (0.21) 799.96 (0.21)
Remark 5.
Recall that for Model $1$, all entries of the discriminant direction $\beta$ and the whitened direction $\zeta$ are non-zero. These entries however can be classified into those representing strong signals vs weak signals based on their magnitudes; see Table 9 in the supplementary. The strong signals appear in the first $10$ entries of $\beta$ (similarly $\zeta$) and the remaining p10p-10 entries of $\beta$ (similarly $\zeta$) correspond to the weak signals. For example, if $\rho=0.5$ then the first $10$ entries of $\beta$ are all equal to $1.98$ and the remaining $p=10$ entries are all equal to 0.02-0.02. From Table 2 we see that $\mathrm{lda}\circ\mathrm{pca}$ kept all of the coordinates corresponding to the strong signals and only added a few coordinates corresponding to the weak signals; this explains the small number of features used in $\mathrm{lda}\circ\mathrm{pca}$. In contrast, the DSDA and SLDA classifiers include a large number of (noisy) features with weak signals.
For Model 2, Table 3 shows that SLDA performs slightly better than $\mathrm{lda}\circ\mathrm{pca}$ and DSDA. However, from Table 4, we see that SLDA also selects a much larger number of features compared to both $\mathrm{lda}\circ\mathrm{pca}$ and DSDA. Recall that, for model $2$, both $\beta$ and $\zeta$ contains exactly $20$ non-zero entries and hence SLDA is selecting a large number of extraneous, non-informative features; a similar, albeit much less severe, phenomenon is observed for DSDA. Table 3 and Table 4 shows that $\mathrm{lda}\circ\mathrm{pca}$, CAT-LDA, and AdaLDA have comparable accuracy with a similar number of selected features. The NSC rule once again has the largest mis-classification error due to it ignoring the correlation structure in $\Sigma$. Finally we see that the rotation pre-processing step described in Hao et al. (2015) lead to a substantial loss in accuracy for the DSDA and AdaLDA classifiers; to understand why this happens, we need to extend the theoretical analysis in Hao et al. (2015) (which assume that $\Sigma$ and $\mathcal{U}_{\mathrm{rot},p}$ are known) to the setting where $\Sigma$ and $\mathcal{U}_{\mathrm{rot},p}$ have to be estimated. We leave this investigation for future work.
Finally for Model 3, Table 5 and Table 6 show that $\mathrm{lda}\circ\mathrm{pca}$ has both the highest accuracy as well as the smallest number of selected features among all the considered classifiers; DSDA has a slightly worse accuracy and also selected a much larger number of features, when compared to $\mathrm{lda}\circ\mathrm{pca}$. SLDA now has the worst accuracy and also selects almost all $p=800$ features, and this is a consequence of $\Sigma$ being a dense matrix. The mis-classification rate for AdaLDA and AdaLDA(rot) are also quite large, and we surmise that this is due to the numerical instability when solving the linear programming problem in AdaLDA. Indeed, the condition numbers for $\Sigma$ can be quite large; see Table 12 in the supplementary for summary statistics of these condition numbers using the same $200$ Monte Carlo replicates as that for generating Table 5. Following the suggestion in Cai & Zhang (2019), we replace the sample covariance matrix $\hat{\Sigma}$ used in the optimization problem for AdaLDA and AdaLDA(rot) (see Eq. (8) and Eq. (9) in Cai & Zhang (2019)) with $\tilde{\Sigma}=\hat{\Sigma}+\sqrt{n^{-1}\ln{p}}\>\mathcal{I}_{p}$; the resulting classifiers are denoted as AdaLDA(reg) and AdaLDA(rot + reg), respectively. Table 5 however shows that using $\tilde{\Sigma}$ only leads to a minimal increase in accuracy. Finally we observe that the rotation pre-processing step once again leads to a substantial loss in accuracy for the DSDA and AdaLDA classifiers.
Remark 6.
While our theoretical results primarily focus on fixed sparsity (see Remark 1 for more on growing sparsity), we deliberately include dense examples with $s_{0}=p$, such as Model 1 and Model 3, to demonstrate robustness beyond the sparse regime. In the dense models, full recovery of $\zeta$ or $\beta$ is not anticipated but $\mathrm{lda}\circ\mathrm{pca}$ still delivers strong classification performance due to effective decorrelation. In particular, for Model 1, our method’s performance is very close to the oracle across varying $\rho$.
Table 3: Mis-classification rate ($\%$) with standard deviations ($\%$) in parentheses for block diagonal setting (model 2), based on $200$ independent Monte Carlo replicates.
$\rho$ 0.50 0.60 0.70 0.80 0.90
Oracle 5.30 (1.71) 5.21 (1.52) 5.34 (1.64) 5.48 (1.62) 5.40 (1.74)
$\mathrm{lda}\circ\mathrm{pca}$ 9.08 (2.71) 8.61 (2.59) 8.30 (2.55) 8.26 (2.45) 8.12 (2.58)
CAT-LDA 10.08 (3.00) 9.36 (2.57) 9.31 (2.50) 9.25 (2.59) 9.21 (2.89)
DSDA 9.63 (2.61) 9.62 (2.64) 9.50 (2.66) 9.58 (2.67) 9.47 (2.42)
DSDA(rot) 20.64 (4.16) 20.52 (3.99) 20.75 (4.18) 21.02 (4.44) 20.90 (4.40)
AdaLDA 12.70 (2.99) 13.06 (3.39) 13.88 (3.45) 13.75 (3.48) 14.68 (3.86)
AdaLDA(rot) 23.91 (3.79) 23.81 (3.72) 24.33 (4.10) 23.89 (3.92) 24.12 (3.94)
SLDA 6.99 (2.89) 6.66 (1.92) 6.71 (2.17) 6.62 (1.97) 6.56 (1.90)
NSC 25.14 (3.11) 26.47 (3.31) 28.61 (3.35) 29.59 (3.24) 30.52 (3.82)
Table 4: Average number of nonzero coefficients with standard deviations in parentheses for block diagonal setting (model 2), based on $200$ independent Monte Carlo replicates.
$\rho$ 0.50 0.60 0.70 0.80 0.90
$\mathrm{lda}\circ\mathrm{pca}$ 20.19 (4.87) 20.48 (5.06) 20.48 (4.88) 20.96 (4.96) 20.58 (4.44)
CAT-LDA 20.02 (4.95) 19.86 (4.31) 20.05 (4.62) 20.32 (4.21) 20.33 (4.54)
DSDA 50.28 (26.61) 50.74 (28.92) 47.48 (25.77) 43.66 (22.42) 45.15 (23.78)
DSDA(rot) 51.02 (39.23) 46.13 (29.09) 51.93 (36.08) 45.56 (31.15) 49.62 (33.50)
AdaLDA 19.45 (2.37) 18.53 (2.23) 17.93 (2.27) 17.40 (2.20) 16.85 (2.07)
AdaLDA(rot) 14.40 (3.14) 14.81 (3.53) 15.16 (3.40) 15.69 (3.99) 14.97 (3.84)
SLDA 405.98 (390.84) 320.17 (380.33) 210.99 (336.12) 242.19 (352.81) 172.00 (309.62)
Table 5: Mis-classification rate ($\%$) with standard deviations ($\%$) in parentheses for random correlation setting (model 3), based on $200$ independent Monte Carlo replicates.
Method $\mathcal{U}(-1,1)$ $\mathcal{N}(0,1)$ $\mathbf{t}_{5}$
$\mathrm{lda}\circ\mathrm{pca}$ 5.07 (2.40) 12.39 (4.17) 13.72 (5.00)
CAT-LDA 11.93 (4.04) 23.67 (6.13) 25.68 (6.82)
DSDA 7.32 (3.05) 16.92 (5.21) 17.90 (5.49)
DSDA(rot) 18.51 (6.37) 31.81 (7.11) 33.65 (7.28)
AdaLDA 17.94 (6.03) 44.55 (7.12) 47.89 (5.10)
AdaLDA(rot) 23.00 (7.49) 39.29 (6.11) 42.52 (6.28)
AdaLDA(reg) 29.48 (7.58) 47.49 (4.52) 48.96 (4.04)
AdaLDA(reg+rot) 21.23 (7.43) 38.93 (6.60) 42.52 (6.28)
SLDA 50.10 (4.40) 49.98 (4.25) 50.09 (3.70)
NSC 30.30 (6.28) 44.55 (5.16) 46.70 (4.49)
Table 6: Model size with standard deviations in parentheses for random correlation setting (model 3), based on $200$ independent Monte Carlo replicates.
Method $\mathcal{U}(-1,1)$ $\mathcal{N}(0,1)$ $\mathbf{t}_{5}$
$\mathrm{lda}\circ\mathrm{pca}$ 11.93 (4.10) 11.48 (3.88) 11.37 (3.92)
CAT-LDA 21.19 (6.51) 20.64 (6.88) 20.75 (6.90)
DSDA 66.00 (31.92) 65.63 (34.45) 64.91 (30.51)
DSDA(rot) 51.08 (34.82) 71.94 (41.64) 77.56 (49.29)
AdaLDA 15.99 (2.32) 10.46 (1.97) 8.51 (2.22)
AdaLDA(rot) 14.62 (4.61) 19.08 (3.93) 16.26 (3.43)
AdaLDA(reg) 11.38 (1.77) 8.50 (2.00) 7.12 (1.96)
AdaLDA(reg+rot) 9.86 (4.71) 15.99 (3.63) 16.26 (3.43)
SLDA 799.51 (1.31) 800.00 (0.07) 800.00 (0.00)
NSC 12.21 (8.47) 55.38 (131.65) 102.89 (192.83)
4.2 Real data analysis
We now assess the performance of $\mathrm{lda}\circ\mathrm{pca}$ on two gene expression data sets for leukaemia and lung cancer. The leukaemia dataset (Golub et al., 1999) includes $p=7128$ gene measurements on $72$ patients with either acute lymphoblastic leukemia (ALL) or acute myeloid leukemia (AML) and we wish to classify patients into either ALL or AML based on their gene expressions. The training set consists of the gene expressions for $n_{1}=27$ patients with ALL and $n_{2}=11$ patients with AML while the test set contains the gene expression data for $20$ patients with ALL and $14$ patients with AML. The lung cancer data was originally analyzed in Gordon et al. (2002) and we use a version of the data wherein the predictor variables with low variances are removed; see Pun & Hadimaja (2021). The resulting data set contains tumor tissues with $p=1577$ features collected from patients with adenocarcinoma (AD) or malignant pleural mesothelioma (MPM). According to Gordon et al. (2002), MPM is highly lethal but distinguishing between MPM and AD is quite challenging in both clinical and pathological settings. The training set consists of $n_{1}=120$ gene expressions for patients with AD and $n_{2}=25$ gene expressions for patients with MPM, and the test set consists of gene expressions for $30$ patients with AD and $6$ patients with MPM. Table 7 and Table 8 presents the classification accuracy and number of selected features for various classifiers when applied to the leukemia and lung cancer data, respectively. The hyperparameters for these classifiers are selected using leave-one-out CV. Table 7 and 8 indicate that the performance of $\mathrm{lda}\circ\mathrm{pca}$ is competitive with existing state-of-the-art classifiers while also operating on a substantially lower-dimensional representation.
Table 7: Mis-classification rate and model size of various methods for the leukaemia data
Method Training Error Test Error Model Size
$\mathrm{lda}\circ\mathrm{pca}$ 0/38 1/34 12
DSDA 0/38 1/34 36
AdaLDA (reg) 0/38 1/34 18
NSC 1/38 3/34 24
Table 8: Mis-classification rate and model size of various methods for the lung cancer data
Method Training Error Test Error Model Size
$\mathrm{lda}\circ\mathrm{pca}$ 0/145 0/36 20
DSDA 1/145 0/36 25
AdaLDA (reg) 0/145 0/36 388
NSC 1/145 0/36 1206
5 Extensions of $\mathrm{lda}\circ\mathrm{pca}$
5.1 Multi-class classification
Suppose we are given training data from $K\geq 3$ classes where the feature vectors for each class are iid samples from a $p$-dimensional multivariate normal distribution, i.e.,
$X_{i1},\dots,X_{in_{i}}\sim\mathcal{N}_{p}(\mu_{i},\Sigma),\qquad i\in[K].$
Here $n_{i}$ denote the number of training data points from class $i\in[K]$. The testing data point $\mathbf{Z}$ is drawn from a mixture of $K$ multivariate normal distributions namely, $\mathbf{Z}\sim\sum_{i=1}^{K}\pi_{i}\mathcal{N}_{p}(\mu_{i},\Sigma)$ with $\pi_{i}\geq 0$ and $\sum_{i=1}^{K}\pi_{i}=1$. Note that, for ease of exposition, the $\{X_{ij}\}$ are assumed to be multivariate normals but the subsequent results also hold when $\{X_{ij}\}$ are elliptically distributed as in Section 7.3.
Now define, for $2\leq i\leq K$, the whitened direction $\zeta^{(i)}$ and the whitened indices $\mathcal{S}_{i}$ via
$\zeta^{(i)}=\mathcal{W}(\mu_{i}-\mu_{1}),\qquad\mathcal{S}_{i}=\{j:\zeta^{(i)}_{j}\neq 0\}.$
Define the global whitened set as $\mathcal{S}_{\zeta}=\mathcal{S}_{2}\cup\mathcal{S}_{3}\cup\dots\cup\mathcal{S}_{K}$. The indices in $\mathcal{S}_{\zeta}$ are the important variables for feature selection. The extension of $\mathrm{lda}\circ\mathrm{pca}$ to $K\geq 3$ classes is then given in Algorithm 3 below, and theoretical results are presented in the supplementary.
Input: $\bar{X}_{i},\>i\in[K]$, $\hat{\Sigma}$ and the test sample $\mathbf{Z}$
Output: $\hat{\Upsilon}_{\mathrm{lda}\circ\mathrm{pca}}(\mathbf{Z})$
Algorithm
 // Step 1: Perform PCA on the feature vectors for the training data (standard PCA approach). Extract the $d$ largest principal components to obtain $\hat{\mathcal{W}}$.
 // Step 2: For $2\leq i\leq K$, let $\tilde{X}_{i}=\hat{\mathcal{W}}\bar{X}_{i}$ and $\hat{\zeta}^{(i)}=\tilde{X}_{i}-\tilde{X}_{1}$. Also let $\hat{\mathcal{S}}_{i}$ be the set of indices corresponding to the $s_{i}$ largest elements of $\hat{\zeta}^{(i)}$ in absolute values. The value of $s_{i}$ is, in general, a user-specified or tuning parameter. Nevertheless, under certain conditions, we can also estimate $s_{i}$; see Eq.˜8.51 in the supplementary.
 // Step 3: Given the test data point $\mathbf{Z}$, let $\tilde{\mathbf{Z}}=\hat{\mathcal{W}}\mathbf{Z}$ and denote by $\hat{\zeta}^{(i)}_{\hat{\mathcal{S}}_{i}}$ the vector obtained from $\hat{\zeta}^{(i)}$ by keeping only those coordinates belonging to $\hat{\mathcal{S}}_{i}$.
 // Step 4: Set $\tilde{D}_{1}=0$ and calculate, for $2\leq i\leq K$, the discriminant score for class $i$ relative to class 11 as
$\tilde{D}_{i}=\bigl[\tilde{\mathbf{Z}}-\bigl(\tfrac{\tilde{X}_{i}+\tilde{X}_{1}}{2}\bigr)\bigr]_{\hat{\mathcal{S}}_{i}}^{\top}\>\hat{\zeta}^{(i)}_{\hat{\mathcal{S}}_{i}}+\ln\frac{n_{i}}{n_{1}}.$ (5.1)
 // Step 5: Assign the label of $\mathbf{Z}$ to the class that maximizes the discriminant score, i.e.,
$\hat{\Upsilon}_{\mathrm{lda}\circ\mathrm{pca}}(\mathbf{Z})=\operatorname*{arg\,max}_{i\in[K]}\tilde{D}_{i}.$ (5.2)
Algorithm 3 $K$-classes $\mathrm{lda}\circ\mathrm{pca}$ decision rule
5.2 Heterogeneous covariance matrices
Despite the simplicity and popularity of regularized or sparse LDA for high-dimensional data, the assumption of equal covariances is not always tenable in practice. More specifically, suppose we are given a $p$-variate random vector $\mathbf{Z}$ drawn from a mixture $\pi_{1}\mathcal{N}_{p}(\mu_{1},\Sigma_{1})+(1-\pi_{1})\mathcal{N}_{p}(\mu_{2},\Sigma_{2})$ with $\Sigma_{1}$ possibly distinct from $\Sigma_{2}$. The Bayes classifier is then the QDA rule given by
$\Upsilon_{F}(\mathbf{Z})=\begin{cases}1&\text{if $\bigl(\mathbf{Z}-\mu_{1}\bigr)^{\top}\Sigma_{1}^{-1}\bigl(\mathbf{Z}-\mu_{1}\bigr)-\bigl(\mathbf{Z}-\mu_{2}\bigr)^{\top}\Sigma_{2}^{-1}\bigl(\mathbf{Z}-\mu_{2}\bigr)\leq\kappa$}\\ 2&\text{otherwise}\end{cases}$ (5.3)
where $\kappa=2\ln\tfrac{\pi_{1}}{1-\pi_{1}}-\ln{\tfrac{|\Sigma_{1}|}{|\Sigma_{2}|}}$. If $\Sigma_{1}=\Sigma_{2}$, then Eq.˜5.3 reduces to Eq.˜1.1. We now discuss how the results in Section˜3 can be extended to quadratic discriminant analysis (QDA) with PCA. Firstly we assume that $\Sigma_{1}$ and $\Sigma_{2}$ both have spiked covariance structures as specified below.
Assumption 5.
Let $\bm{u}_{i1},\dots,\bm{u}_{id_{i}},$ for $i=1,2$ be orthonormal vectors in $\mathbb{R}^{p}$ and assume that the covariance matrix $\Sigma_{i}$ for the $p$-variate normal distributions $\mathcal{N}_{p}(\mu_{i},\Sigma_{i})$ is of the form
$\Sigma_{i}$ $=\>\sum_{k=1}^{d_{i}}\lambda_{ik}\mathbf{u}_{ik}\mathbf{u}_{ik}^{\top}+\sigma_{i}^{2}\mathcal{I}_{p}=\mathcal{U}_{i}\Lambda_{i}\mathcal{U}_{i}^{\top}+\sigma_{i}^{2}\mathcal{I}_{p},\quad i=1,2.$ (5.4)
Here $\Lambda_{i}=\mathrm{diag}(\lambda_{ik})$ is a $d_{i}\times d_{i}$ diagonal matrix and $\mathcal{U}_{i}$ is a $p\times d_{i}$ matrix with orthonormal columns. We assume implicitly that $\lambda_{i1}\geq\cdots\geq\lambda_{id}>0$, $\sigma_{i}>0$ and $d_{i}\ll p,\>i=1,2$.
Note that a recent line of research on QDA for high-dimensional classification is based on the assumption that $\Sigma_{2}^{-1}-\Sigma_{1}^{-1}$ is sparse; see e.g., Li & Shao (2015); Jiang et al. (2018); Cai & Zhang (2021). In contrast, ˜5 do not enforce any sparsity assumption and also allows for $\Sigma_{1}$ and $\Sigma_{2}$ to have different spiked eigenvalues and eigenvectors; the latter is is a generalization of the common principal components (CPC) assumption in Zhu (2006); Pepler et al. (2017); Flury (1988) where the leading principal components are the same for both $\Sigma_{1}$ and $\Sigma_{2}$.
Under ˜5, the whitening transformation $\mathcal{W}_{i}=\Sigma_{i}^{-1/2}$ for $i=1,2$ is of the form in Eq.˜2.7 and thus a suitable estimate for $\mathcal{W}_{i}$ is given by Eq.˜3.3. Let $\zeta_{i}=\mathcal{W}_{i}\mu_{i}$ for $i=1,2,$ and denote the whitened index sets by $\mathcal{A}_{i}=\{j:\zeta_{ij}\neq 0\}$ for $i=1,2$. Let $\mathcal{A}_{0}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$ and note that the elements in $\mathcal{A}_{0}$ are the signal coordinates for the QDA rule (after the PCA step). Eq.˜5.3 can be written as
$\Upsilon(\mathbf{Z}):=\begin{cases}1&\text{if $\|[\mathcal{W}_{1}\bigl(\mathbf{Z}-\mu_{1}\bigr)]_{\mathcal{A}_{0}}\|^{2}-\bigl\|[\mathcal{W}_{2}\bigl(\mathbf{Z}-\mu_{2}\bigr)]_{\mathcal{A}_{0}}\|^{2}\leq\kappa$ }\\ 2&\text{otherwise}\end{cases}$ (5.5)
A plugin decision rule is then obtained by replacing $\mu_{i},\Sigma_{i}$ and $\mathcal{A}_{0}$ with their estimates $\bar{X}_{i},\hat{\Sigma}_{i}$ and $\hat{\mathcal{A}}_{0}$. Note that the intercept $\kappa$ is non-trivial to estimate in the high-dimensional setting as it involves the log-determinant $\ln{\tfrac{|\Sigma_{1}|}{|\Sigma_{2}|}}$; see Cai et al. (2015) for further details. For our paper we employ the data-driven approach of Jiang et al. (2018) which circumvents the need to estimate the determinants of $\Sigma_{1}$ and $\Sigma_{2}$. The full details of $\mathrm{qda}\circ\mathrm{pca}$ are specified in Algorithm 4 below, and theoretical results for $\mathrm{qda}\circ\mathrm{pca}$ are provided in the supplementary.
Input: $\bar{X}_{1}$, $\bar{X}_{2}$, $\hat{\Sigma}_{1}$, $\hat{\Sigma}_{2}$ and a test data point $\mathbf{Z}$
10pt
Output: $\hat{\Upsilon}_{\mathrm{qda}\circ\mathrm{pca}}(\mathbf{Z})$
10pt
Algorithm
 // Step 1: Perform PCA on the feature vectors for the training data (standard PCA approach). Extract, for $i=1,2,$ the $d_{i}$ largest principal components $\hat{\mathcal{U}}_{i}$ and compute $\hat{\mathcal{W}}_{i}$ as in Eq.˜$3.$
 // Step 2: For $i=1,2,$, set $\hat{\zeta}_{i}=\hat{\mathcal{W}}_{i}\bar{X}_{i}$ and form the indices set $\hat{\mathcal{A}}_{i}$ by selecting the $s_{i}$ largest (in magnitude) coordinates of $\hat{\zeta}_{i}$.
 // Step 3:Let $\hat{\mathcal{A}}_{0}=\hat{\mathcal{A}}_{1}\cup\hat{\mathcal{A}}_{2}$ and define for any $\bm{x}\in\mathbb{R}^{p}$
$\mathbf{Q}(x\mid\{\bar{X}_{1},\bar{X}_{2},\hat{\mathcal{W}}_{1},\hat{\mathcal{W}}_{2},\hat{\mathcal{A}}_{0}\})=\bigl\|[\hat{\mathcal{W}}_{1}(\bm{x}-\bar{X}_{1})]_{\hat{\mathcal{A}}_{0}}\bigr\|^{2}-\bigl\|[\hat{\mathcal{W}}_{2}(\bm{x}-\bar{X}_{2})]_{\hat{\mathcal{A}}_{0}}\bigr\|^{2}.$
 // Step 4:Find $\hat{\kappa}$ to minimize the empirical 0-11 loss of the decision rule induced by $\mathbf{Q}$, i.e.,
$\hat{\kappa}=\operatorname*{arg\,min}_{\eta\in\mathbb{R}}\frac{1}{n}\sum_{i=1}^{2}\sum_{j=1}^{n_{i}}\mathbbm{1}(\Upsilon(X_{ij}\mid\{\bar{X}_{1},\bar{X}_{2},\hat{\mathcal{W}}_{1},\hat{\mathcal{W}}_{2},\hat{\mathcal{A}}_{0},\eta_{\mathrm{thresh}}\})\neq i)$ (5.6)
where, for any $\eta_{\mathrm{thresh}}\in\mathbb{R}$, we define
$\Upsilon(X_{ij}\>|\,\{\bar{X}_{1},\bar{X}_{2},\hat{\mathcal{W}}_{1},\hat{\mathcal{W}}_{2},\hat{\mathcal{A}}_{0},\eta_{\mathrm{thresh}}\})=\begin{cases}1&\text{if $\mathbf{Q}(X_{ij}\mid\{\bar{X}_{1},\bar{X}_{2},\hat{\mathcal{W}}_{1},\hat{\mathcal{W}}_{2},\hat{\mathcal{A}}_{0}\})\leq\eta_{\mathrm{thresh}}$}\\ 2&\text{otherwise}\end{cases}$
 // Step 5: Given a test data point $\mathbf{Z}$, return the decision rule
$\hat{\Upsilon}_{\mathrm{qda}\circ\mathrm{pca}}(\mathbf{Z})=\begin{cases}1&\text{if $\mathbf{Q}(\mathbf{Z}\>|\,\{\bar{X}_{1},\bar{X}_{2},\hat{\mathcal{W}}_{1},\hat{\mathcal{W}}_{2},\hat{\mathcal{A}}_{0}\})\leq\hat{\kappa}$}\\ 2&\text{otherwise}\end{cases}$ (5.7)
Algorithm 4 $\mathrm{qda}\circ\mathrm{pca}$ decision rule
As indicated by Jiang et al. (2018), $\eta_{\mathrm{thresh}}$ is selected by minimizing the in-sample misclassification error based on $\bar{X}{1}$, $\bar{X}{2}$, $\hat{\Sigma}{1}$, and $\hat{\Sigma}{2}$. Under ˜5, the grid search can be narrowed to a neighborhood around $2\log\left(\frac{|\hat{\mathcal{W}}{2}|}{|\hat{\mathcal{W}}{1}|}\right)-2\log\left(\frac{n_{1}}{n_{2}}\right)$. This is justified by the improved stability of the log-determinant ratio in the spiked setting, where bulk eigenvalue $\sigma_{i}^{2}$ is consistently estimated via the pooled sample covariance and the leading spikes remain well-separated (from the bulks).
6 Discussion
In this paper we addressed the classification problem for high-dimensional data by analyzing the prototypical $\mathrm{lda}\circ\mathrm{pca}$ classifier that first transforms the feature vectors using a whitening transformation, then performs feature selection on the whitened data, and finally applies LDA in the dimensionally reduced space. We show that, under a spiked eigenvalue structure for $\Sigma$, the mis-classification error rate for $\mathrm{lda}\circ\mathrm{pca}$ is asymptotically Bayes optimal whenever $n\rightarrow\infty$ and $n^{-1}\ln p\rightarrow 0$. While the Bayes consistency of $\mathrm{lda}\circ\mathrm{pca}$ is similar to that of classifiers based on estimating the discriminant direction $\beta$, the underlying assumptions and motivations for our results are substantially different. Indeed, the focus on PCA and the whitening matrix leads to the natural assumption that $\Sigma$ has spiked/diverging eigenvalues while earlier results that focus on estimation of $\beta$ had generally assumed that $\Sigma$ is sparse or that the eigenvalues of $\Sigma$ are bounded. Numerical experiments indicate that $\mathrm{lda}\circ\mathrm{pca}$ is competitive with existing state-of-the-art classifiers while operating on a substantially lower-dimensional representation. This behavior persists even when the underlying sparsity levels in the raw and whitened feature spaces are matched; see more details in Section˜7.2.
We now mention two interesting issues for future research. The first is to extend the theoretical results in this paper for combining LDA with PCA to other, possibly non-linear, dimension reduction techniques such as (classical) multidimensional scaling, kernel PCA, and Laplacian eigenmaps (Belkin & Niyogi, 2003), followed by learning a classifier in the dimensionally reduced space. The second issue concerns the spiked covariance structure in ˜1. In particular, while ˜1 is widely used in the literature, see e.g., Hao et al. (2015); Cai et al. (2013); Johnstone (2001); Birnbaum et al. (2012), its assumption on the non-spiked eigenvalues and eigenvectors might be somewhat restrictive. We can consider relaxing ˜1 by assuming that $\Sigma$ arises from an approximate factor model (Fan et al., 2013) or that $\Sigma$ can be decomposed into a low-rank plus sparse matrix structure (Agarwal et al., 2011). We surmise that, due to the focus on the whitening matrix $\Sigma^{-1/2}$, theoretical analysis of $\mathrm{lda}\circ\mathrm{pca}$ under these more general covariance structure will also leads to interesting technical developments; e.g., while perturbation results for $\hat{\Sigma}^{-1}-\Sigma^{-1}$ given $\hat{\Sigma}-\Sigma$ are well-studied, much less is known about perturbation bounds for $\hat{\Sigma}^{-1/2}-\Sigma^{-1/2}$ given $\hat{\Sigma}-\Sigma$.